# weighted-amenability and percolation

### v1

@@ 3. $\sqrt{\mathbf{w}}$-biasing
We apply this theory to an explicit random walk whose bias accounts for the weight function $\mathbf{w}$. Consider a $\Gamma$ closed subgroup of $\operatorname{Aut}(G)$ that acts transitively on $G$ and let $\mathbf{w}^{o}(\cdot)$ be the induced relative weight function on the vertices as in (1) with respect to some reference point $o\in V$. Besides edge weights being constant $1$, the most natural network to be associated with $(G,\mathbf{w}^{o})$ is the one whose stationary measure is $\mathbf{w}^{o}$. This is the network where the weight of edge $(x,y)$ will be $\widetilde{\mathbf{w}}^{o}(x,y):=\sqrt{\mathbf{w}^{o}(x)\mathbf{w}^{o}(y)}$ (or any multiple of it by a positive constant). Indeed, letting the transition probability from $x$ to $y$ be $p^{(\mathbf{w})}(x,y)\propto\widetilde{\mathbf{w}}^{o}(x,y)$, we have reversibility:
$\mathbf{w}^{o}(x)p^{(\mathbf{w})}(x,y)$ $=\mathbf{w}^{o}(x)\frac{\widetilde{\mathbf{w}}^{o}(x,y)}{\sum_{z\sim x}\widetilde{\mathbf{w}}^{o}(x,z)}=\mathbf{w}^{o}(x)\frac{\widetilde{\mathbf{w}}^{o}(x,y)}{\frac{\mathbf{w}^{o}(x)}{\mathbf{w}^{o}(y)}\sum_{z\sim y}\widetilde{\mathbf{w}}^{o}(y,z)}$
$=\mathbf{w}^{o}(y)\frac{\widetilde{\mathbf{w}}^{o}(x,y)}{\sum_{z\sim y}\widetilde{\mathbf{w}}^{o}(y,z)}=\mathbf{w}^{o}(y)p^{(\mathbf{w})}(y,x).$
@@
$c(\partial_{E}F)$ $=\sum_{x\in F\,,\,y\in\partial_{V}F\atop x\sim y}\mathbf{w}^{o}(y)p^{(\mathbf{w})}(y,x)\geqslant\frac{1}{\sqrt{d}D^{\mathbf{w}}}\sum_{y\in\partial_{V}F}\mathbf{w}^{o}(y)=\frac{1}{\sqrt{d}D^{\mathbf{w}}}\cdot\mathbf{w}^{o}(\partial_{V}F).$
$\blacksquare$
@@
$=\sum_{\delta>1}\left(\delta+1-2\sqrt{\delta}\right)N_{\delta}>0.$
$\blacksquare$
@@
Let $\Gamma$ be a closed subgroup of $\operatorname{Aut}(G)$ that acts transitively on $G$ and let $\mathbf{w}$ be the induced relative weight function as in (1). Suppose $X^{(\mathbf{w})}_{n}$ is the $\sqrt{\mathbf{w}}$-biased random walk on $G$, then $Y_{n}:=\log\left(\mathbf{w}^{o}\left(X^{\mathbf{w}}_{n}\right)\right)$ is a symmetric random walk on a countable subset of $\mathbb{R}$.
@@
$=\operatorname{\mathds{P}}_{o}\left(\mathbf{w}^{o}\left(X^{(\mathbf{w})}_{1}\right)=\frac{1}{\delta}\right).$
$\blacksquare$
@@
$\rho(G)$ $=\lim\sup_{n\to\infty}p_{2n}(o,o)^{\frac{1}{2n}}\leqslant\sup_{n\to\infty}p_{2n}^{\mathbf{w}}(o,o)^{\frac{1}{2n}}=\rho^{\mathbf{w}}(G).$
$\blacksquare$
@@ 4. Level-amenability and hyperfiniteness
Definition 4.1.
(Hyperfiniteness) Let $\Gamma$ be a closed subgroup of $\operatorname{Aut}(G)$ that acts transitively on $G$ and $\mathbf{w}$ be the induced relative weight function as in (1). Suppose $\omega$ is an invariant random subgraph of $G$. We say that $\omega$ is hyperfinite if there is a family of stochastically increasing $\Gamma$-invariant bond (or site) percolation processes $\{\omega_{n}\}_{n\in\mathbb{N}}$ on $G$ such that
@@
    the increasing union of configurations of $\omega_{n}$ is equal to $\omega$,
@@
We may assume that configurations $\omega_{n}$ sampled from such a family of percolation $\mathbf{P}_{n}$ are monotonically coupled (otherwise switch to $\widehat{\omega}_{m}:=(\cup_{k=m}^{\infty}\omega_{n_{k}}^{c})^{c}$ for a suitable subsequence $n_{k}$), and thus we will write $\omega_{n}\nearrow\omega$ if $\omega_{n}\subseteq\omega_{n+1}$ for all $n\in\mathbb{N}$ and $\cup_{n}\omega_{n}=\omega$.
The above definition can be viewed as an adaptation of the notions of hyperfiniteness for graphings (or for measurable equivalence relations) and unimodular random graphs [AL07]. In these cases hyperfiniteness was also known to be equivalent to the respective notion of amenability [CFW81, Kai97].
@@
    In Definition 1.3 one can replace ``$\left|F\right|<\infty"$ by ``$\mathbf{w}^{o}(F)<\infty"$, i.e.,
@@
Notice that $\mathbf{w}^{o}(\cdot)$ restricted to a slice takes only finitely many values, thus each slice is a quasi-transitive unimodular graph (see [Tan19, Proposition 4.3] for a proof). We are now ready to prove Theorem 1.4.
Proof.
@@
where $\widehat{D}(\cdot,\cdot)$ denotes the the weighted degree as in (5) and $d$ is the degree in $G$. The conclusion now follows, from Theorem 2.16. $\blacksquare$
@@
As in the proof of Theorem 1.4, we define Bernoulli$(p)$ level-percolation of a nonunimodular graph $G$ by assigning i.i.d. Bernoulli($p$) random variables to each level of $G$ and retaining only those edges whose endpoints are in levels with label 1. In particular, for any $p\in(0,1)$ such a process splits $G$ into a union of slices. The following easy corollary will be handy in a later construction.
@@
Let $\Gamma$ be a closed subgroup of $\operatorname{Aut}(G)$ that acts transitively on $G$ and $\mathbf{w}$ be the induced relative weight function as in (1). If $G$ is $\mathbf{w}$-nonamenable, then for any $p\in(0,1)$ there is a connected component of Bernoulli$(p)$ level-percolation that is a nonamenable quasi-transitive unimodular graph.
@@ 5. Invariant random spanning trees and forests
[…]
First, we recall that Free Minimal Spanning Forest on the graph $G$ is a random subforest, denoted as $\mathrm{FMSF}(G)$, that is constructed as follows:
@@
Given a relative weight function $\mathbf{w}^{o}:V\to\mathbb{R}^{+}$ with respect to some reference vertex $o\in V$ the Free $\mathbf{w}$-Maximal Spanning Forest $\mathrm{FMaxSF}_{\mathbf{w}}(G)$ is constructed in a similar fashion:
@@
First we would like to highlight some of the results from [CTT22].
@@
Let $\Gamma$ be a closed subgroup of $\operatorname{Aut}(G)$ that acts transitively on $G$ and $\mathbf{w}$ be the induced relative weight function as in (1). Suppose $p\in[0,1]$ is such that there are infinitely many heavy clusters $\operatorname{\mathds{P}}_{p}$-a.s. Then for $\operatorname{\mathds{P}}_{p}$-a.e. configuration $\omega$, the random forest $\mathrm{FMaxSF}_{\mathbf{w}}(\omega)$ almost surely contains infinitely many trees with infinitely many $\mathbf{w}$-nonvanishing ends, none of which is isolated.
In the earlier version of the manuscript [CTT22] this statement had a slightly stronger assumption, namely that $p\in(p_{h},p_{u})$, but the proof remains the same after its relaxation.
Theorem 5.2 ([CTT22, Corollary 5.7]).
Let $\Gamma$ be a closed subgroup of $\operatorname{Aut}(G)$ that acts transitively on $G$ and $\mathbf{w}$ be the induced relative weight function as in (1). Let $\mathbf{P}$ be a $\Gamma$-invariant percolation such that a.s.  there a cluster with $\geqslant 3$ $\mathbf{w}$-nonvanishing ends. Then the cluster graphing $\mathcal{G}^{\mathrm{cl}}$ with the measure induced by $\mathbf{P}$ is $\mu$-nonamenable. In fact, it contains a $\mu$-nonamenable Borel subforest.
These results naturally motivate a question of characterization of $\mathbf{w}$-amenability in terms of the number of $\mathbf{w}$-nonvanishing ends in $\Gamma$-invariant random subforests.
@@
Let $\Gamma$ be a closed subgroup of $\operatorname{Aut}(G)$ that acts transitively on $G$ and $\mathbf{w}$ be the induced relative weight function as in (1). Suppose $\mathbf{P}$ is a $\Gamma$-invariant random spanning forest. Then $\mathbf{P}$-a.s.  for every heavy cluster $T$ the following are equivalent:
@@
    $p_{h}(T)<1$,
@@
    $T$ has $\geqslant 3$ $\mathbf{w}$-nonvanishing ends.
@@
    $T$ has $\geqslant 3$ ends with heavy geodesics.
Proof.
The equivalence of (1), (3), and (4) was proven in [TTD25], as in Theorem 2.21, in the context of acyclic Borel graphs. One can derive our version of the statements from it by considering cluster graphings induced by the percolation process $\mathbf{P}$ and applying Theorem 2.21 together with Observation 2.23. Below we provide the equivalence of (2) to these, and an alternative proof to (1$\Leftrightarrow$(3).
(1$\Rightarrow$(2): Assume $p_{h}(T)=1$. Associate an independent Uniform $[0,1]$ random variable $U_{e}$ to every edge $e\in\varphi$ and for $\varepsilon\in[0,1]$ let $T_{1-\varepsilon}$ consist of those edges that have $U_{e}\leqslant 1-\varepsilon$. For every $\varepsilon>0$ all clusters in $T_{1-\varepsilon}$ are light and $T_{1-\varepsilon}\nearrow T$ as $\varepsilon\searrow 0$. By Lemma 4.2.(2), $T$ is hyperfinite.
(2) $\Rightarrow$(3): The assumption $p_{h}(T)<1$ and $p_{u}(T)=1$ (since $T$ is a tree) imply that there is some $p<1$ such that Bernoulli$(p)$ percolation on $T$ has infinitely many heavy clusters. By Lemma 2.20 each such cluster contains a $\mathbf{w}$-nonvanishing end. Since $T$ is a tree, the ends of different clusters must correspond to distinct ends in $T$.
(3) $\Rightarrow$ (1): Call a vertex $x\in T$ a $\mathbf{w}$-trifurcation if its removal disconnects $T$ into at least three components each of which contains a $\mathbf{w}$-nonvanishing end. Since $T$ has $\geqslant 3$ $\mathbf{w}$-nonvanishing ends such vertices exist. We will first show that almost surely every tree $T$ contains infinitely many $\mathbf{w}$-trifurcation vertices of the same weight. Suppose $T$ that does not have this property. Notice that then there must exist $k(T)\in\mathbb{N}$ and an infinitely increasing set of levels in which T has exactly $k$ $\mathbf{w}$-trifurcation vertices (call such levels good), otherwise choosing one $\mathbf{w}$-trifurcation vertex is the highest such level and letting every vertex in $T$ send unit mass to it violates TMTP. Moreover, we may assume that there exists $r\geqslant 1$ such that there are infinitely many good levels with weights in $\{r^{m}\}_{m\in\mathbb{Z}}$ with respect to some reference vertex $o$. (Otherwise define a mass transport where every vertex $x$ sends mass 1 to each $\mathbf{w}$-trifurcation in a level $L$ that is good and has weight in $\{\mathbf{w}^{o}(x)r^{m}\}_{m\in\mathbb{Z}}$, whenever $L$ has maximal and/or minimal weight of all such levels. If there existed such an $L$ then the expected mass received would be infinite, while the expected mass sent out would be at most $2k$.) Finally, consider the following mass transport scheme: each $\mathbf{w}$-trifurcation vertex in $T$ from a good level sends a unit mass to every $\mathbf{w}$-trifurcation vertex in $T$ from every good level above it whose weight differs by $r^{m}$ for some $m\in\mathbb{Z}$. Clearly, the mass sent out is infinite, while the expected received mass is finite as there is an exponential decay in the weights below.
Assume now that every level in $T$ contains none or infinitely many $\mathbf{w}$-trifurcation vertices. For each level $L$ with infinitely many $\mathbf{w}$-trifurcation one can construct (using new edges that are given by a function of the original edges in $F$) an invariant forest on the $\mathbf{w}$-trifurcation vertices such that every tree has infinitely many ends. Namely, add an edge between $\mathbf{w}$-trifurcation vertices $x$ and $y$ in $L$ if $y$ is the closest $\mathbf{w}$-trifurcation vertex to $x$ in a component it disconnects, and label this new edge with the distance of $x$ and $y$ in $F$. This graph has infinitely many ends. Take a unimodular random subforest $F^{\prime}$ where some trees have infinitely many ends, as in [AL07, Theorem 8.13]. The expected degree of a vertex conditioned in being in such a tree is greater than 2, so there is $k$ large enough such that there is a unimodular random subforest consisting of trees $F^{\prime}_{k}\subseteq F^{\prime}$ that contains only edges with labels $\leqslant k$ and has infinitely many ends, and hence it is not hyperfinite by [AL07, Corollary 8.10]. $\blacksquare$
@@
One cannot replace $\mathbf{w}$-nonvanishing ends by heavy ends in the statement of Theorem 5.3 and, as a consequence, in Theorems 1.6 and 5.5. Indeed, consider $T_{1,2}$ from Example 2.7. Such a tree is hyperfinite, has a unique nonvanishing end, but every end in this tree is heavy. We delve deeper in the relations between various notions of ends at the end of this section.
@@
    For any $\Gamma$-invariant random spanning forest $\mathbf{F}$ of $G$, the $\mathbf{F}$-configuration $\varphi$ contains only trees with $\leqslant 2$ $\mathbf{w}$-nonvanishing ends a.s.
@@
    For any $\Gamma$-invariant random spanning forest $\mathbf{F}$ of $G$, the $\mathbf{F}$-configuration $\varphi$ satisfies $p_{h}(\varphi)=1$ a.s.
@@
(1$\Rightarrow$(4). Suppose there is a tree in $\varphi$ with at least $3$ $\mathbf{w}$-nonvanishing ends with positive probability. By Theorem 2.21 the corresponding cluster graphing is $\mu$-nonamenable. Observation 2.23(c) now yields nonamenability of $\Gamma$. The $\mathbf{w}$-nonamenability of $G$ now follows from Theorem 2.3.
(4$\Rightarrow$(1). Suppose $G$ is $\mathbf{w}$-nonamenable by Corollary 4.5 Bernoulli$(p)$ level-percolation on $G$ has a connected component that induces a nonamenable quasi-transitive graph. By [BLS99, Theorem 3.10] adopted to the quasi-transitive setting we have that such a component admits an invariant spanning forest with positive Cheeger constant and, thus, infinitely many ends. Since a slice contains only finitely many levels each these ends is $\mathbf{w}$-nonvanishing.
(4$\Leftrightarrow$(5). Follows from Theorem 5.3. $\blacksquare$
@@
(1$\Leftrightarrow$(2) is part of Theorem 5.5.
(1$\Rightarrow$(3) The forward direction was shown in Theorem 1.6, as a $\Gamma$-invariant random spanning tree $T$ with $\leqslant 2$ $\mathbf{w}$-nonvanishing ends has $p_{h}(T)=1$ by Theorem 5.3.
(1$\Leftarrow$(3) The other direction follows from hyperfiniteness and is based on the first part of the argument in the proof of [LP16, Theorem 8.21]. Assign to each edge $e\in E(G)$ an independent Uniform$[0,1]$ random variable $U_{e}$. Given a random connected subgraph $\omega$ let $\omega_{1-1/n}$ be its subgraph that contains all of the edges with $U_{e}\leqslant 1-1/n$. Since $p_{h}(\omega)=1$ all clusters in $\omega_{1-1/n}$ are light. For every vertex $x\in V$ define $W(x)$ to be the closest vertex (in $G$-distance) in $\omega$ to $x$, if there are several then choose one uniformly at random. Let $\xi_{n}$ consist of all edges $(x,y)$ such that $W(x)$ and $W(y)$ are in the same cluster in $\omega_{1-1/n}$. By the TMTP, for all $n$ every cluster in $\xi_{n}$ has to be light and $\xi_{n}\nearrow G$. Hence $G$ is hyperfinite which is equivalent to $\mathbf{w}$-amenability by Theorem 1.4.
(1$\Leftrightarrow$(4) This was proved by Benjamini, Lyons, Peres and Schramm, as in Theorem 1.5. $\blacksquare$
@@
It follows directly from considering the cluster graphings induced by the percolation process $\mathbf{P}$ and applying Theorem 2.21 and Observation 2.23 . $\blacksquare$
Lemmas 5.3 and 5.6 naturally motivate the question about the relations between various notions of ends introduced in Subsection 2.4 at least in the context of invariant random subforests.
@@
[…]

### v2

@@ 3. $\sqrt{\mathbf{w}}$-biasing
We apply this theory to an explicit random walk whose bias accounts for the weight function $\mathbf{w}$. Consider a $\Gamma$ closed subgroup of $\operatorname{Aut}(G)$ that acts transitively on $G$ and let $\mathbf{w}^{o}(\cdot)$ be the induced relative weight function on the vertices as in (1) with respect to some reference point $o\in V$. Besides edge weights being constant $1$, the most natural network to be associated with $(G,\mathbf{w}^{o})$ is the one whose stationary measure is $\mathbf{w}^{o}$. This is the network where the weight of edge $(x,y)$ are $\widetilde{\mathbf{w}}^{o}(x,y):=\sqrt{\mathbf{w}^{o}(x)\mathbf{w}^{o}(y)}$ (or any multiple of it by a positive constant). Indeed, letting the transition probability from $x$ to $y$ be $p^{(\mathbf{w})}(x,y)\propto\widetilde{\mathbf{w}}^{o}(x,y)$, we have reversibility:
$\mathbf{w}^{o}(x)p^{(\mathbf{w})}(x,y)$ $=\mathbf{w}^{o}(x)\frac{\sqrt{\mathbf{w}^{o}(x)\mathbf{w}^{o}(y)}}{\sum_{z\sim x}\sqrt{\mathbf{w}^{o}(x)\mathbf{w}^{o}(z)}}=\frac{\sqrt{\mathbf{w}^{o}(x)\mathbf{w}^{o}(y)}}{\sum_{z\sim x}\sqrt{\frac{\mathbf{w}^{o}(z)}{\mathbf{w}^{o}(x)}}}$
$=\frac{\sqrt{\mathbf{w}^{o}(x)\mathbf{w}^{o}(y)}}{\sum_{z\sim y}\sqrt{\frac{\mathbf{w}^{o}(z)}{\mathbf{w}^{o}(y)}}}=\mathbf{w}^{o}(y)\frac{\sqrt{\mathbf{w}^{o}(x)\mathbf{w}^{o}(y)}}{\sum_{z\sim y}\sqrt{\mathbf{w}^{o}(y)\mathbf{w}^{o}(z)}}=\mathbf{w}^{o}(y)p^{(\mathbf{w})}(y,x),$
where in the first equality in the second line we used transitivity of $G$ and $\Gamma$-invariance of $\mathbf{w}$.
@@
$c(\partial_{E}F)$ $=\sum_{x\in F\,,\,y\in\partial_{V}F\atop x\sim y}\mathbf{w}^{o}(y)p^{(\mathbf{w})}(y,x)\geqslant\frac{1}{\sqrt{d}D^{\mathbf{w}}}\sum_{y\in\partial_{V}F}\mathbf{w}^{o}(y)=\frac{1}{\sqrt{d}D^{\mathbf{w}}}\cdot\mathbf{w}^{o}(\partial_{V}F).$ $\blacksquare$
@@
$=\sum_{\delta>1}\left(\delta+1-2\sqrt{\delta}\right)N_{\delta}>0.$ $\blacksquare$
@@
Let $\Gamma$ be a closed subgroup of $\operatorname{Aut}(G)$ that acts transitively on $G$ and let $\mathbf{w}$ be the induced relative weight function as in (1). Suppose $X^{(\mathbf{w})}_{n}$ is the $\sqrt{\mathbf{w}}$-biased random walk on $G$. Then $Y_{n}:=\log\left(\mathbf{w}^{o}\left(X^{\mathbf{w}}_{n}\right)\right)$ is a symmetric random walk on a countable subset of $\mathbb{R}$.
@@
$=\operatorname{\mathds{P}}_{o}\left(\mathbf{w}^{o}\left(X^{(\mathbf{w})}_{1}\right)=\frac{1}{\delta}\right).$ $\blacksquare$
@@
$\rho(G)$ $=\lim\sup_{n\to\infty}p_{2n}(o,o)^{\frac{1}{2n}}\leqslant\sup_{n\to\infty}p_{2n}^{\mathbf{w}}(o,o)^{\frac{1}{2n}}=\rho^{\mathbf{w}}(G).$ $\blacksquare$
@@ 4. Level-amenability and hyperfiniteness
Definition 4.1 (Hyperfiniteness).
Let $\Gamma$ be a closed subgroup of $\operatorname{Aut}(G)$ that acts transitively on $G$ and $\mathbf{w}$ be the induced relative weight function as in (1). Suppose $\omega$ is an invariant random subgraph of $G$. We say that $\omega$ is hyperfinite if there is an increasing family random subgraphs $\omega_{1}\subseteq\omega_{2}\subseteq\ldots\subseteq\omega_{n}\subseteq\ldots\subseteq\omega$ such that
@@
    the joint law of $(\omega,\{\omega_{n}\}_{n\in\mathbb{N}})$ is $\Gamma$-invariant,
  2. (2)
    $\cup\omega_{n}=\omega$,
@@
We will write $\omega_{n}\nearrow\omega$ if $\omega_{n}\subseteq\omega_{n+1}$ for all $n\in\mathbb{N}$ and $\cup_{n}\omega_{n}=\omega$.
The above definition can be viewed as an adaptation of the notions of $\mu$-hyperfiniteness for graphings (or for measurable equivalence relations) and unimodular random graphs [AL07].
@@
    In Definition 1.3 one can replace $\left|F\right|<\infty$" by $\mathbf{w}^{o}(F)<\infty$", i.e.,
@@
This notion of hyperfiniteness for a random subgraph of $G$ is connected to hyperfiniteness from measure group theory via the cluster graphing construction (see Subsection 2.5).
Lemma 4.3.
Let $\Gamma$ be a closed subgroup of $\operatorname{Aut}(G)$ that acts transitively on $G$ and $\mathbf{w}$ be the induced relative weight function as in (1). Suppose $\omega$ is an invariant random subgraph of $G$. Then $\omega$ is hyperfinite if and only if $\omega$ admits a cluster graphing $\mathcal{G}^{\mathrm{cl}}$ on a standard probability space $(Y,\nu)$ that is $\nu$-hyperfinite.
Proof.
$\Rightarrow:$ Let $\omega_{n}$ be a sequence of random subgraphs that witnesses hyperfiniteness of $\omega.$ Let $\mathbf{P}$ denote the joint law of $(\omega,\{\omega_{n}\}_{n\in\mathbb{N}})$ and consider the joint cluster graphing, that is the same construction as in the summary at the end Subsection 2.5, but started with an ergodic free pmp action of $\Gamma$ on the standard probability space $(X,\mu)$ that factors onto $\left((2^{E})^{\mathbb{N}},\mathbf{P}\right)$. This yields an increasing family of cluster graphings $\mathcal{G}^{\mathrm{cl}}_{\omega_{n}}$ on $(Y,\nu)$ with finite connected components on a $\nu$-conull set (because each of these connected components is isomorphic to a cluster in the corresponding configuration $\omega_{n}$), and $\mathcal{G}^{\mathrm{cl}}_{\omega}=\bigcup_{n}\mathcal{G}^{\mathrm{cl}}_{\omega_{n}}$. Thus $\mathcal{G}^{\mathrm{cl}}_{\omega}$ is $\nu$-hyperfinite.
$\Leftarrow:$ Suppose the cluster graphing $\mathcal{G}^{\mathrm{cl}}_{\omega}$ of $\omega$ is an increasing union of Borel graphs $\{\mathcal{G}_{n}\}_{n\in\mathbb{N}}$ whose connected components are finite. By Observation 2.23 each $\mathcal{G}^{\mathrm{cl}}_{\omega}$-connected component of $\widetilde{x}=[(x,o)]_{\Gamma}\in Y$ is isomorphic to the connected component of $o$ in $\pi(x)$, thus the images under this isomorphism of the $\mathcal{G}_{n}$ yield an increasing exhausting family of invariant random subgraphs of $\omega$. In particular, $\omega$ is hyperfinite. $\blacksquare$
@@
Notice that $\mathbf{w}^{o}(\cdot)$ restricted to a slice takes only finitely many values, thus each slice is a quasi-transitive unimodular graph (see [Tan19, Proposition 4.3] for a proof).
Observation 4.5.
Let $\Gamma$ be a closed subgroup of $\operatorname{Aut}(G)$ that acts transitively on $G$ and $\mathbf{w}$ be the induced relative weight function as in (1). Define the level graph $\mathcal{L}(G)$ as the graph on the set of levels of $G$ with edges induced by the adjacency in $G$. Then $\mathcal{L}$ is the Cayley graph of a finitely generated Abelian group. In particular, it is hyperfinite.
Given an invariant random partition of the set of levels of $G$, we define a level-percolation on $G$ is defined on by retaining only those edges whose endpoints are in levels in the same class in the partition. A hyperfinite exhaustion of the level graph $\mathcal{L}(G)$ defines an increasing family of invariant random partitions such that each class is a slice and their union is $G$. We are now ready to prove Theorem 1.4.
Proof of Theorem 1.4.
@@
[…]
@@
The following easy corollary will be handy in a later construction.
@@
Let $\Gamma$ be a closed subgroup of $\operatorname{Aut}(G)$ that acts transitively on $G$ and $\mathbf{w}$ be the induced relative weight function as in (1). If $G$ is $\mathbf{w}$-nonamenable, then there is an invariant level-percolation on $G$ that a.s. contains a connected component that is a nonamenable quasi-transitive unimodular graph.
@@ 5. Invariant random spanning trees and forests
An essential tool for many percolation results in the unimodular setting is [BLPS99b, Proposition 7.2]. It connects the number of ends of an invariant random spanning forest $\mathcal{F}$ to $p_{c}(\mathcal{F})$, and to the average degree in the infinite components of $\mathcal{F}$. In addition to their applications in percolation theory, invariant spanning forests also played an important role in measured group theory, especially in the pmp setting. Famously, this machinery enabled Gaboriau and Lyons to prove a positive answer to a measure theoretic version of the Day–von Neumann conjecture [GL09]. As we highlighted in Subsection 2.5, extensions of this theory to the mcp setting have received an increasing amount of attention in recent years. In this section, we consider invariant random spanning trees and forests in the context of $\mathbf{w}$-amenability and nonunimodular transitive graphs.
First, we recall that the Free Minimal Spanning Forest on the graph $G$ is a random subforest that is constructed as follows:
@@
In [CTT22] the authors introduced a weighted analog of the Free Minimal Spanning Forest (FMSF) called the Free $\mathbf{w}$-Maximal Spanning Forest, denoted as $\mathrm{FMaxSF}_{\mathbf{w}}$, and showed that it naturally extends many desired properties of FMSF to the nonunimodular setting. Given a relative weight function $\mathbf{w}^{o}:V\to\mathbb{R}^{+}$ with respect to some reference vertex $o\in V$ the Free $\mathbf{w}$-Maximal Spanning Forest $\mathrm{FMaxSF}_{\mathbf{w}}(G)$ is constructed in a similar fashion:
@@
First we would like to highlight some of the results from [CTT22]. We present the first one for Bernoulli(p)(p) percolation, however, the original result holds more generally for insertion and deletion tolerant $\Gamma$-invariant percolation processes see [CTT22, Theorem 4.16 and Remark 4.17].
Theorem 5.1 ([CTT22]).
Let $\Gamma$ be a closed subgroup of $\operatorname{Aut}(G)$ that acts transitively on $G$ and $\mathbf{w}$ be the induced relative weight function as in (1). Suppose $p\in[0,1]$ is such that there are infinitely many heavy clusters $\operatorname{\mathds{P}}_{p}$-a.s. Then for $\operatorname{\mathds{P}}_{p}$-a.e. configuration $\omega$, for every heavy cluster $C\subseteq\omega$, the random forest $\mathrm{FMaxSF}_{\mathbf{w}}(C)\subseteq\mathrm{FMaxSF}_{\mathbf{w}}(\omega)$ a.s. contains a tree $T\subseteq C$ with infinitely many $\mathbf{w}$-nonvanishing ends, none of which is isolated.
@@
Let $\Gamma$ be a closed subgroup of $\operatorname{Aut}(G)$ that acts transitively on $G$ and $\mathbf{w}$ be the induced relative weight function as in (1). Let $\mathbf{P}$ be a $\Gamma$-invariant percolation such that with positive probability there a cluster with $\geqslant 3$ $\mathbf{w}$-nonvanishing ends. Then the cluster graphing $\mathcal{G}^{\mathrm{cl}}$ with the measure induced by $\mathbf{P}$ is $\mu$-nonamenable. In fact, it contains a $\mu$-nonamenable Borel subforest.
The following theorem is largely an adaptation of the results due to Adams and Lyons [AL91] (see also [JKL02, Section 3.6]) and [TTD25], as in Theorem 2.21, to our setting. Our main addition to the statement is the third condition that is similar to the $p_{c}$ condition from [BLPS99b, Proposition 7.2].
@@
Let $\Gamma$ be a closed subgroup of $\operatorname{Aut}(G)$ that acts transitively on $G$ and $\mathbf{w}$ be the induced relative weight function as in (1). Suppose $\mathcal{F}$ is a $\Gamma$-invariant random spanning subforest of $G$. Then almost surely the following are equivalent:
@@
    On a possibly larger probability space there is an invariant way to select an end in every infinite tree in $\mathcal{F}$.
@@
    For every heavy tree $T\in\mathcal{F}$ we have $p_{h}(T)=1$ a.s.
@@
    Every heavy tree $T\in\mathcal{F}$ has $\leqslant 2$ $\mathbf{w}$-nonvanishing ends.
  5. (5)
    Every heavy tree $T\in\mathcal{F}$ has $\leqslant 2$ ends with heavy geodesics.
Remark 5.4.
Without enlarging the probability space (2) can be changed to the existence of an invariant selection at most two ends or finitely many vertices in every tree in $\mathcal{F}$.
Proof of Theorem 5.3.
[…]
Consider a cluster graphing $\mathcal{G}^{\mathrm{cl}}$ of $\mathcal{F}$. Then by Lemma 4.3 $\mathcal{F}$ is hyperfinite if and only if $\mathcal{G}^{\mathrm{cl}}$ is $\nu$-hyperfinite. Now, by [JKL02, Lemma 2.21], in a.e. $\mathcal{G}^{\mathrm{cl}}$-connected component we may select $\leqslant 2$-ends or finitely many points in a Borel way. By Observation 2.23, $\mathcal{G}^{\mathrm{cl}}$ components are isomorphic to corresponding clusters of $\mathcal{F}$. Hence the selection in $\mathcal{G}^{\mathrm{cl}}$ induces a selection in the clusters of $\mathcal{F}$. Clearly, by the TMTP a.s. in every heavy tree in $\mathcal{F}$ we selected $1$ or $2$ ends. In case of one we are done, while in case of two choosing between selected ends at random yields the desired selection.
[…]
@@
One cannot replace $\mathbf{w}$-nonvanishing ends by heavy ends in the statement of Theorem 5.3 and, as a consequence, in Theorems 1.6 and 5.6. Indeed, consider $T_{1,2}$ from Example 2.7. Such a tree is hyperfinite, has a unique nonvanishing end, but every end in this tree is heavy. We delve deeper in the relations between various notions of ends at the end of this section.
@@
    For any $\Gamma$-invariant random spanning forest $\mathcal{F}$ of $G$, the $\mathcal{F}$-configuration $\varphi$ contains only trees with $\leqslant 2$ $\mathbf{w}$-nonvanishing ends a.s.
@@
    For any $\Gamma$-invariant random spanning forest $\mathcal{F}$ of $G$, the $\mathcal{F}$-configuration $\varphi$ satisfies $p_{h}(\varphi)=1$ a.s.
@@
[…]
@@
[…]
@@
It follows directly from considering the cluster graphing induced by the percolation process $\mathbf{P}$ and applying Theorem 2.21 and Observation 2.23. $\blacksquare$
Theorem 5.3 and Lemma 5.7 naturally motivate the question about the relations between various notions of ends introduced in Subsection 2.4 at least in the context of invariant random subforests.
@@
Case 2: two nonvanishing ends. Suppose a heavy component-tree contains exactly two nonvanishing ends. Then by a simple application of TMTP all vanishing ends are light, and thus have only light geodesics. Indeed, let every vertex in such a tree send a unit mass to the closest point on the unique bi-infinite geodesic path that connects two nonvanishing ends. Similarly, the $\limsup$’s of weights along geodesic paths to nonvanishing ends must be equal to each other and greater than zero. Assuming the values of $\limsup$’s are different, orient the edges on the bi-infinite geodesic from the end with the lower value towards the one with the higher value. There must be the first vertex (in this orientation) whose weight is larger than the average of the $\limsup$’s. Since this point was chosen invariantly, this violates TMTP. Finally, assume both $\limsup$’s are zero then there are finitely many vertices on the bi-infinite path of maximal weight, this again contradicts TMTP.